# Decomposing Representation Space into Interpretable Subspaces with Unsupervised Learning

### v1

@@ E.1 Building Activation Database
As we mentioned in main paper, we apply InversionView to subspaces and simply do vector search to find preimages. Here are the details: (1) The text data includes 2000 random sequences in MiniPile, 1000 IOI prompts and 1000 Greater-than prompts. We made IOI and Greater-than prompts using the same procedure as we did for clean inputs in patching experiments. We run GPT-2 Small on these text, producing 1178970 activations for each activation site (i.e., post MLP residual stream at layer 4, 6, 8, 9). Each activation has a corresponding token position from which it is collected. We maintain their correspondence such that we can always retrieve the input tokens that result in certain activation (i.e., all tokens in its previous context). (2) For a subspace partition trained for an activation site, i.e., orthogonal matrix $\mathbf{R}$ and configuration $c$, we make $S$ (number of subspaces) databases by using FAISS [Douze et al., 2024] index. Activations in whole space are converted into subspace activations (multiplied by $\mathbf{R}$ and split according to $c$) and are added to each database respectively. Therefore, we make separate databases for each layer (activation site) and each subspace. (3) Then for any interested subspace activation (query activation), we search for other activations in corresponding subspace database whose cosine similarity is above the threshold. We then show the inputs corresponding to query activation and searching results.
@@ E.2 Input Attribution Details
As language model’s representations are contextualized, sometimes activations encode information far from the token position it corresponds to, i.e., long-range dependency. This makes interpretation difficult, as we need to search commonality in a large context. We thus make use of input attribution method to highlight important tokens. We use integrated gradients with 10 interpolations. Specifically, given an query subspace activation and a sequence of tokens, we first obtain the initial residual stream activations from these tokens, and multiply them with $\alpha=0.1,0.2,\dots,1.0$ to create 10 interpolations. We continue the forward pass to get subspace activations. Importantly, we freeze all attention weights to the values they take when running on clean inputs ($\alpha=1.0$) during the forward pass. We finally compute the cosine similarity between resulting subspace activation and the given query subspace activation. The gradient (not via attention weights) of the cosine similarity with respect to each interpolations are averaged and multiplied with residual stream activations to obtain the final token-level attribution score. The scores are normalized by the sum of scores of all tokens in the same input sequence before showing in the APP. So the scores sum to 1. But they usually spread somewhat evenly across many tokens, resulting in invisible highlights, so we also provide options in the APP to scale these scores together. Because it requires time to compute attribution scores, we pre-compute them for 50 random query activations and show them in a separate page in the APP.
The reason for fixing attention weights is that, we empirically find it produces more plausible attribution. The idea behind it is to emphasize what is moved by the attention circuit, while ignoring how this circuit is formed. For example, for a induction head on input “A B … A”, it would only highlight “B” though all 3 tokens play a role. The same intuition is also used in other works [Ameisen et al., 2025]. Importantly, we find the effectiveness of our attribution method is limited, as the scores are usually too concentrated at the most recent token, readers should be cautious when using them. This also highlights importance of better input attribution, as in larger and more capable models long-range dependency might be more prevalent.
@@
Figure 22: Subspace 7 at $\boldsymbol{h}^{4,post}$: Preimages of the residual stream activation corresponding to two tokens adjacent to the one in Fig. 2(c) (“under”, “different”), and two tokens in very different positions (at position 565 and 384) in the same subspace after the 5th layers in GPT-2 Small. Interpretation of encoded information: (a) current position is around 118. (b) current position is around 119. (c) current position is around 565. (d) current position is around 384. In this subspace, one can increase threshold for cosine similarity and get enough samples of very similar position.
@@ E.3 More examples and Detailed Findings
Fig. 20 - 23 correspond to 4 subfigures in the main paper Fig. 2, in each of them we fix the subspace and vary inputs to observe the consistency.
@@
Figure 25: Subspace 4 at $\boldsymbol{h}^{4,post}$: Preimages of the residual stream activation. Different from other preimage figures, we show input attribution scores here, because the encoded information depends on long-range context, we need these hints to help us find commonality. The blue color represent the attribution score (see Sec. E.2 for details), they are amplified by a factor of 20 for visibility. The token where the activation is taken from is marked by bold and underscored font. Encoded information is: (a) the current token is repeated multiple times. (b) The current token is repeated multiple times. Note that we check the entire previous context of the query activation, and there are indeed more “water” tokens there. We don’t show it due to space limit. (c) The current token “above” (if we only consider similarity $>$ 0.85) (d) The current token “ac”. We checked the entire context and found the token “ac” is not repeated. Overall, we found the following pattern: when the current token is relatively less frequent and is repeated, this subspace clusters together activations corresponding to different tokens, otherwise it clusters activations according to current token. The signal might be important to activate induction heads in later layers.
@@
We also find the remaining two subspaces 9 and 10 hard to explain. For subspace 9, we find that, somewhat surprisingly, almost all activations are in a similar direction, as shown in Fig. 27. We also find the norm of activations in this subspace varies, the main variance in this subspace is from this direction. If this is a single feature, what is the pattern that affect the magnitude in this direction? However, we do not have very confident conclusion for this question. We find that the norm is almost always small if the current token is in the middle of a multi-token word (e.g., Fig. 27(b)’s bottom row) and very big 2000$\approx$ 2000≈ 2000 for “end of sequence” (EOS) token. One point we can believe is that, the information encoded in this direction is quite independent with the information in other subspaces. Interestingly, subspace 9 in $\boldsymbol{h}^{8,post}$ also shows this pattern, i.e., all activations point to similar direction and big norm for EOS. But subspace in that layer is the one with the highest effect in test 3 described in Sec. 5.1: it has an un-normalized effect of 2.31 ($d_{9}=32$), the second and third largest is 0.78 ($d_{2}=96$) and 0.30. This means such an subspace indeed encodes certain meaningful information. In IOI domain it is the position of the subject (outputs from S-Inhibition heads), but it might be part of a more general mechanism, like a signal to not attend to certain part of context. That explains why it is so hard to interpret.
Come back to layer 4, for subspace 10, we also see many activations in the same direction, but not as often as subspace 9. For both 9 and 10, we do not have a clear conclusion for the information encoded by them. But we believe they are worth more extensive study, as it might reveal some unexpected features used by the model.
@@
After exhaustively exploring subspaces in the residual stream after layer 4 (the 5th layer), we move on to take a brief look at subspaces in later layers. We show some preimages in Fig. 28 and Fig. 29, which correspond to two subspaces with the highest patching effect in quantitative evaluation. The APP includes results for all 4 layers (4, 6, 8, 9) used in quantitative evaluation, we encourage readers to check our APP.
To check interpretability of subspaces n later layers, we also try to interpret every subspace of layer 9, and summarize our findings as follows:
@@
In Fig. 4(a), we see that there are two subspaces, 2 and 6, playing important roles for parametric knowledge routing. We apply the same qualitative analysis to see what is the information encoded in them, and if they differ in some way. Like in quantitative tests, we test and remove entities that model does not have knowledge for. We randomly sample 250 entity pairs from them, each is used to construct a 2-shot example. We then run Qwen2.5-1.5B on these examples, together with 2000 text sequences from MiniPile, and collect activations to build the database. Fig. 30 shows some examples. As we can see, they indeed differ. Though they both contain retrieved parametric knowledge, subspace 2 encodes the knowledge about the current n-gram, while subspace 6 encodes knowledge that does not depend on current token. We therefore have the following hypothesis: when enough tokens are observed and the name is identifiable, MLP retrieves knowledge and places it into subspace 2, this knowledge is then emitted by attention heads to many other tokens’ residual stream (or in other words, attended by other tokens), this information arrives at certain subspace similar to subspace 6, but at later layers. This behavior is probably distributed and repeated across multiple middle layers, so subspace 6 in this layer receives information from certain subspaces in earlier layer that resemble subspace 2 of this layer. So the information transfer is not done all at once. To verify this hypothesis, more rigorous experiments should be done. We encourage readers to check more examples. Due to its large size and various limits we did not manage to deploy the database for Qwen2.5-1.5B on our web application. But we have released all trained orthogonal matrices together with the code. Readers can follow the instructions in our repository and easily reproduce the database.
Appendix F Additional Discussion
F.1 Why is InversionView More Suitable in Subspaces

### v2

@@ E.1 Building Activation Database
As we mentioned in main paper, we apply InversionView to subspaces and simply do vector search to find preimages. Here are the details: (1) The text data includes 2000 random sequences in MiniPile, 1000 IOI prompts and 1000 Greater-than prompts. We made IOI and Greater-than prompts using the same procedure as we did for clean inputs in patching experiments. We run GPT-2 Small on these text, producing 1178970 activations for each activation site (i.e., post MLP residual stream at layer 4, 6, 8, 9). Each activation has a corresponding token position from which it is collected. We maintain their correspondence such that we can always retrieve the input tokens that result in certain activation (i.e., all tokens in its previous context). (2) For a subspace partition trained for an activation site, i.e., orthogonal matrix $\mathbf{R}$ and configuration $c$, we make $S$ (number of subspaces) databases by using FAISS (Douze et al., 2024) index. Activations in whole space are converted into subspace activations (multiplied by $\mathbf{R}$ and split according to $c$) and are added to each database respectively. Therefore, we make separate databases for each layer (activation site) and each subspace. (3) Then for any interested subspace activation (query activation), we search for other activations in corresponding subspace database whose cosine similarity is above the threshold. We then show the inputs corresponding to query activation and searching results.
@@ E.2 Input Attribution Details
As language model’s representations are contextualized, sometimes activations encode information far from the token position it corresponds to, i.e., long-range dependency. This makes interpretation difficult, as we need to search commonality in a large context. We thus make use of input attribution method to highlight important tokens. We use integrated gradients with 10 interpolations. Specifically, given an query subspace activation and a sequence of tokens, we first obtain the initial residual stream activations from these tokens, and multiply them with $\alpha=0.1,0.2,\dots,1.0$ to create 10 interpolations. We continue the forward pass to get subspace activations. Importantly, we freeze all attention weights to the values they take when running on clean inputs ($\alpha=1.0$) during the forward pass. We finally compute the cosine similarity between resulting subspace activation and the given query subspace activation. The gradient (not via attention weights) of the cosine similarity with respect to each interpolations are averaged and multiplied with residual stream activations to obtain the final token-level attribution score. The scores are normalized by the sum of scores of all tokens in the same input sequence before showing in the web application. So the scores sum to 1. But they usually spread somewhat evenly across many tokens, resulting in invisible highlights, so we also provide options in the web application to scale these scores together. Because it requires time to compute attribution scores, we pre-compute them for 50 random query activations and show them in a separate page in the application.
The reason for fixing attention weights is that, we empirically find it produces more plausible attribution. The idea behind it is to emphasize what is moved by the attention circuit, while ignoring how this circuit is formed. For example, for a induction head on input “A B … A”, it would only highlight “B” though all 3 tokens play a role. The same intuition is also used in other works (Ameisen et al., 2025). Importantly, we find the effectiveness of our attribution method is limited, as the scores are usually too concentrated at the most recent token, readers should be cautious when using them. This also highlights importance of better input attribution, as in larger and more capable models long-range dependency might be more prevalent.
@@
Figure 23: Subspace 7 at $\boldsymbol{h}^{4,post}$: Preimages of the residual stream activation corresponding to two tokens adjacent to the one in Fig. 3(c) (“under”, “different”), and two tokens in very different positions (at position 565 and 384) in the same subspace after the 5th layers in GPT-2 Small. Interpretation of encoded information: (a) current position is around 117. (b) current position is around 119. (c) current position is around 565. (d) current position is around 384. For this subspace, one can increase the threshold and obtain samples of very similar positions as shown in Figure 31
@@ E.3 More examples and Detailed Findings
Fig. 21 - 24 correspond to 4 subfigures in the main paper Fig. 3, in each of them we fix the subspace and vary inputs to observe the consistency.
@@
Figure 26: Subspace 4 at $\boldsymbol{h}^{4,post}$: Preimages of the residual stream activation. Different from other preimage figures, we show input attribution scores here, because the encoded information depends on long-range context, we need these hints to help us find commonality. The blue color represent the attribution score (see Sec. E.2 for details), they are amplified by a factor of 20 for visibility. The token where the activation is taken from is marked by bold and underscored font. Encoded information is: (a) the current token is repeated multiple times. (b) The current token is repeated multiple times. Note that we check the entire previous context of the query activation, and there are indeed more “water” tokens there. We don’t show it due to space limit. (c) The current token “above” (if we only consider similarity >> 0.85) (d) The current token “ac”. We checked the entire context and found the token “ac” is not repeated. Overall, we found the following pattern: when the current token is relatively less frequent and is repeated, this subspace clusters together activations corresponding to different tokens, otherwise it clusters activations according to current token. The signal might be important to activate induction heads in later layers.
@@
We also find the remaining two subspaces 9 and 10 hard to explain. For subspace 9, we find that, somewhat surprisingly, almost all activations are in a similar direction, as shown in Fig. 28. We also find the norm of activations in this subspace varies, the main variance in this subspace is from this direction. If this is a single feature, what is the pattern that affect the magnitude in this direction? However, we do not have very confident conclusion for this question. We find that the norm is almost always small if the current token is in the middle of a multi-token word (e.g., Fig. 28(b)’s bottom row) and very big $\approx 2000$ for “end of sequence” (EOS) token. One point we can believe is that, the information encoded in this direction is quite independent with the information in other subspaces. Interestingly, subspace 9 in $\boldsymbol{h}^{8,post}$ also shows this pattern, i.e., all activations point to similar direction and big norm for EOS. But subspace in that layer is the one with the highest effect in test 3 described in Sec. 5.1: it has an un-normalized effect of 2.31 ($d_{9}=32$), the second and third largest is 0.78 ($d_{2}=96$) and 0.30. This means such an subspace indeed encodes certain meaningful information. In IOI domain it is the position of the subject (outputs from S-Inhibition heads), but it might be part of a more general mechanism, like a signal to not attend to certain part of context. That explains why it is so hard to interpret.
Coming back to layer 4, for subspace 10, we also see many activations in the same direction, but not as often as subspace 9. For both 9 and 10, we do not have a clear conclusion for the information encoded by them. But we believe they are worth more extensive study, as it might reveal some unexpected features used by the model.
@@
Figure 31: Subspace 7 at $\boldsymbol{h}^{4,post}$: Same examples as Figure 23(d), but the threshold for cosine similarity is increased to 0.99, we obtain samples of very similar positions. Remaining error could be because the optimization is not optimal, or the model has forgotten the information a bit at this layer.
@@
After exhaustively exploring subspaces in the residual stream after layer 4 (the 5th layer), we move on to take a brief look at subspaces in later layers. We show some preimages in Fig. 29 and Fig. 30, which correspond to two subspaces with the highest patching effect in quantitative evaluation. The web application includes results for all 4 layers (4, 6, 8, 9) used in quantitative evaluation, we encourage readers to check our web application.
To check interpretability of subspaces in later layers, we also try to interpret every subspace of layer 9, and summarize our findings as follows:
@@
In Fig. 5(a), we see that there are two subspaces, 2 and 6, playing important roles for parametric knowledge routing. We apply the same qualitative analysis to see what is the information encoded in them, and if they differ in some way. Like in quantitative tests, we test and remove entities that model does not have knowledge for. We randomly sample 250 entity pairs from them, each is used to construct a 2-shot example. We then run Qwen2.5-1.5B on these examples, together with 2000 text sequences from MiniPile, and collect activations to build the database. Fig. 32 shows some examples. As we can see, they indeed differ. Though they both contain retrieved parametric knowledge, subspace 2 encodes the knowledge about the current n-gram, while subspace 6 encodes knowledge that does not depend on current token. We therefore have the following hypothesis: when enough tokens are observed and the name is identifiable, MLP retrieves knowledge and places it into subspace 2, this knowledge is then emitted by attention heads to many other tokens’ residual stream (or in other words, attended by other tokens), this information arrives at certain subspace similar to subspace 6, but at later layers. This behavior is probably distributed and repeated across multiple middle layers, so subspace 6 in this layer receives information from certain subspaces in earlier layer that resemble subspace 2 of this layer. So the information transfer is not done all at once. To verify this hypothesis, more rigorous experiments should be done. We encourage readers to check more examples. Due to its large size and various limits we did not manage to deploy the database for Qwen2.5-1.5B on our web application. But we released all trained orthogonal matrices together with the code. Readers can follow the instructions in our repository and easily reproduce the database.
Appendix F Comparison to Sparse Autoencoders (SAEs)
Due to the popularity of SAEs, in this section, we explain the difference and connections between them and our method. Readers who are familiar with SAEs might find this section particularly helpful.
F.1 FAQ
(1) Why not compare with SAEs?
Quantitative evaluation
In previous quantitative evaluations, we patch each subspace and see its average effect on the output. SAEs do not provide such subspaces directly. If we have to make a comparison, we can change the evaluation method to apply an analogous evaluation to SAEs. One way is to patch the basis elements that SAEs provide, the features. This means that we replace a feature’s activation value with its value on the corrupted input. For example, in the clean input on the token after "John", a feature representing *"previous token is John"* is highly activated – while, in the same position in the corrupted input, a feature representing *"previous token is Mary"* is highly activated. If we patch the "prev-Mary" feature, the new activation would say *"previous token is both John and Mary"*. If we patch the "prev-John" feature, the new activation has no information about previous token. In other words, we need to add a feature while removing another one (and it’s hard to select which one since there are multiple active features), which requires understanding of feature dependency that SAEs do not provide ("prev-Mary" and "prev-John" are mutually exclusive). More importantly, patching the "prev-John" feature would not have any real effect on inputs where the names are not John. In other words, conceptually, one would expect such patching to only have effect on a few specific inputs, and only to confuse the model. In contrast, patching subspace can be thought of patching meaningful feature groups. Moreover, applying such patching to SAEs requires running evaluation on each of the features, which is too computationally expensive due to the large amount of features.
Because of these fundamental difference, we do not compare with SAEs in the main paper. Nevertheless, readers might still wonder how would SAEs perform under the same evaluation used in the paper. For this, a better way is to derive subspaces from SAE features. We take this up in Section F.2.
Qualitative evaluation
Our qualitative evaluation has two aspects: (i) interpretability of preimages for subspaces, (ii) consistency of the type of information encoded in a subspace across inputs (thus, interpretability of the overall subspace). Regarding (i), we think that a single SAE feature would have interpretability comparable to a single subspace preimage, as we know that they are interpretable most of the time but also some features do not have clear and easy interpretation. However, regarding (ii), this is not applicable to SAEs, as they provide no clear structure between features; rather, to find encodings of the same type of information requires inspecting each feature individually. In contrast, information in the same subspace found by NDM is of the same "type" across inputs most of the time. Inspecting some of the them can give an idea of the whole feature group, significantly reducing the amount of interpretation work needed.
(2) What about training SAEs within subspaces?
Conceptually, if one trained a top-1 SAE (top-k SAE with k=1) within each subspace, we expect that the features would be the cluster centers in the subspace. Inspecting a feature means to find inputs whose activation in the subspace is close to the feature. This is roughly the same as inspecting preimages in InversionView, with query activations taking the role of the "feature". Thus, InversionView can conceptually be thought of as a "parameter-free" top-1 SAE.
F.2 Quantitative Experiments on GPT2 Test Suite
In this section, we check whether we can obtain some meaningful subspaces from SAE features and evaluate them with our metrics.
Settings
We use the method introduced in Engels et al. (2024) (Sec. 4), where the authors cluster feature vectors and find subspaces. Specifically, the method has three steps:
  1. 1.
    Cluster features by their pairwise cosine similarity. There are two clustering method, (a) Graph. It connects each feature with its top-k neighbors whose similarity is above a threshold, and form clusters by recursively searching for connected features. (b) Spectral. Spectral clustering.
  2. 2.
    For each cluster, run the SAE on model activations. When reconstructing activations, they only use features inside the cluster (i.e., activation value for other features are always zero), we refer them as cluster-specific reconstructions. They remove zero reconstructions from this collection, so that at least one feature in the cluster is activated.
  3. 3.
    Find the multi-dimensional subspaces spanned by these cluster-specific reconstructions. They apply PCA on these reconstructions and inspect the points in the subspaces spanned by pairs of principle components with large eigenvalues.
In our case, we aim to extract subspaces from clusters of features. Thus we slightly change their method for our purpose:
  1. 1.
    Cluster features. Same as above.
  2. 2.
    Sort clusters either by the average cosine similarity between features inside the cluster, or by the cluster size (number of features) in descending order.
    For each cluster (in this order), apply the next three steps. Overall, we iteratively build a matrix $C\in\mathbb{R}^{d\times m}$ of orthonormal where $m$ is the number of dimensions selected so far. The columns of $C$ are orthonormal columns, representing a basis for the subspaces constructed so far. for each cluster, we add another set of columns representing the subspace for that cluster.
    1. (a)
      Obtain cluster-specific reconstructions, using the same method as Step 2 above in the original method. Denote the reconstructions as $X\in\mathbb{R}^{d\times n}$, where $n$ is the number of data points and $d$ is model dimension.
    2. (b)
      Remove variance on previous selected directions (skip if no direction has been selected so far). Denote the directions selected so far as $C\in\mathbb{R}^{d\times m}$, where $m$ is the number of selected directions. We then take
      $\hat{X}\leftarrow X-CC^{T}X$ (13)
      which removes variance on the directions selected so far, or – equivalently – projects on the orthogonal complement of the subspaces constructed so far.
    3. (c)
      Apply SVD on $\hat{X}$. We select all left singular vectors (i.e., the same same as the eigenvectors in PCA) whose corresponding singular values are above a threshold. The threshold is set by multiplying the maximum singular value with a hyperparameter rtol. The selected singular vectors, or directions, form a new subspace corresponding to this feature cluster. They are added as the new columns of $C$.
  3. 3.
    We stop once $C$ becomes a full basis, i.e., $C\in\mathbb{R}^{d\times d}$, or the clusters are exhausted. In the latter case, we supplement $C$ with remaining directions to make it an orthogonal matrix.
In sum, we follow the original method from Engels et al. (2024) as much as we can, while adapting it to construct a full orthogonal decomposition of the space. We test this method on the GPT-2 test suite, with many different hyperparameter configurations. Each configuration is a combination of the following hyperparameters: In step 1, use Graph or Spectral. For the former, there are two hyperparameters: the k used in defining top-k neighbors and the threshold cutoff. For the latter, num specifies how many cluster there should be. In step 2, sort the feature in two ways, avg-sim or size. In Step 5, rtol that filters out unimportant directions. We search k among [2, 3, 4, 5, 6, 8, 16, 32], search cutoff among [0.1, 0.11, 0.13, 0.14, 0.16, 0.18, 0.21, 0.23, 0.26, 0.3, 0.34, 0.38, 0.43, 0.48, 0.5, 0.55, 0.62, 0.7, 0.78, 0.89, 1.0], search num among [10, 25, 50, 100], search rtol among [0.01, 0.03, 0.1]. For graph, the resulting number of clusters varies a lot. We remove those configurations whose number of clusters in any layer is out of the range [5, 200]. This results in 16 configurations for graph. When extracting subspaces, we ignore clusters which contain $\leq 3$ features (as the target information we are testing is not likely to be encoded in 3D or smaller subspaces), and further restrict the number of subspace to be inside [5, 50] in order to speed up the evaluation. Unexpectedly, none of the configuration using graph produced a partition satisfying the constraint. We find that this is mainly because it produces very uneven-sized clusters. There are usually a few big clusters, and many small clusters each of which has a few features. Some configurations that use spectral are end up being filtered out.
Results
The results are shown in Table 8. We can see there are two best-performing configurations whose average Gini coefficient is 0.56; we put the first row in the main paper, Table 1. It shows that extracting subspaces from SAE feature clusters is significantly better than other baseline methods (0.21-0.38 in Table 1), probably producing somewhat meaningful partition. But there is still a significant gap between it and NDM’s result (0.71). In addition, we can see that many configurations are somewhat successful in test 1, e.g., the third to last row (Spectral, num=100, avg-sim, rtol=0.03), but not very successful in other tests.
Cluster config Extract config # subspaces test 1 test 2 test 3 test 4 test 5 Avg
- $d_{s}$ $\text{Var}_{s}$ - $d_{s}$ $\text{Var}_{s}$ - $d_{s}$ $\text{Var}_{s}$ - $d_{s}$ $\text{Var}_{s}$ - $d_{s}$ $\text{Var}_{s}$
Spectral, num=10 avg-sim, rtol=0.1 7,8,9,7 0.60 0.44 0.60 0.61 0.41 0.67 0.73 0.39 0.54 0.74 0.31 0.54 0.71 0.56 0.61 0.56
Spectral, num=10 size, rtol=0.1 5,7,8,8 0.70 0.61 0.35 0.40 0.32 0.56 0.70 0.30 0.49 0.64 0.22 0.44 0.68 0.50 0.46 0.49
Spectral, num=25 avg-sim, rtol=0.03 6,7,7,7 0.78 0.72 0.72 0.42 0.16 0.36 0.48 0.24 0.33 0.35 0.20 0.30 0.48 0.33 0.41 0.42
Spectral, num=25 avg-sim, rtol=0.1 10,12,12,13 0.52 0.60 0.53 0.50 0.26 0.48 0.60 0.43 0.54 0.60 0.30 0.51 0.59 0.36 0.48 0.49
Spectral, num=25 size, rtol=0.1 9,9,8,7 0.81 0.67 0.36 0.53 0.48 0.52 0.55 0.38 0.33 0.49 0.44 0.42 0.59 0.51 0.42 0.50
Spectral, num=50 avg-sim, rtol=0.01 5,6,6,6 0.65 0.67 0.71 0.37 0.15 0.42 0.13 0.11 0.35 0.18 0.14 0.35 0.27 0.15 0.32 0.33
Spectral, num=50 avg-sim, rtol=0.03 7,8,9,8 0.77 0.69 0.73 0.45 0.21 0.41 0.43 0.19 0.36 0.39 0.18 0.34 0.43 0.25 0.46 0.42
Spectral, num=50 avg-sim, rtol=0.1 15,18,21,19 0.85 0.80 0.53 0.55 0.31 0.44 0.52 0.29 0.36 0.46 0.27 0.42 0.49 0.26 0.49 0.47
Spectral, num=50 size, rtol=0.1 7,8,8,7 0.63 0.62 0.68 0.45 0.29 0.35 0.52 0.43 0.49 0.46 0.35 0.38 0.64 0.48 0.50 0.48
Spectral, num=100 avg-sim, rtol=0.01 9,8,8,9 0.70 0.56 0.72 0.28 0.17 0.36 0.21 0.09 0.27 0.19 0.11 0.26 0.31 0.14 0.30 0.31
Spectral, num=100 avg-sim, rtol=0.03 13,12,11,12 0.84 0.70 0.81 0.45 0.21 0.47 0.34 0.16 0.32 0.33 0.18 0.33 0.34 0.18 0.36 0.40
Spectral, num=100 avg-sim, rtol=0.1 24,26,26,25 0.86 0.74 0.54 0.51 0.33 0.47 0.48 0.32 0.43 0.41 0.24 0.42 0.38 0.21 0.41 0.45
Spectral, num=100 size, rtol=0.1 9,12,12,11 0.76 0.74 0.66 0.54 0.49 0.57 0.63 0.59 0.60 0.51 0.33 0.29 0.69 0.59 0.40 0.56
Table 8: Results on GPT-2 Small test suite, higher is better. Refer to Table 1 for explanation of the metrics. We include hyperparameters and the number of subspaces produced for each layer used in the tests.
Appendix G Additional Discussion
G.1 Why is InversionView More Suitable in Subspaces